# Advancing Large Language Models to Capture Varied Speaking Styles and Respond Properly in Spoken Conversations

### v1

@@ Abstract
In spoken dialogue, even if two current turns are the same sentence, their responses might still differ when they are spoken in different styles. The spoken styles, containing paralinguistic and prosodic information, mark the most significant difference between text and speech modality. When using text-only LLMs to model spoken dialogue, text-only LLMs cannot give different responses based on the speaking style of the current turn. In this paper, we focus on enabling LLMs to listen to the speaking styles and respond properly. Our goal is to teach the LLM that "even if the sentences are identical if they are spoken in different styles, their corresponding responses might be different". Since there is no suitable dataset for achieving this goal, we collect a speech-to-speech dataset, StyleTalk, with the following desired characteristics: when two current speeches have the same content but are spoken in different styles, their responses will be different. To teach LLMs to understand and respond properly to the speaking styles, we propose the Spoken-LLM framework that can model the linguistic content and the speaking styles. We train Spoken-LLM using the StyleTalk dataset and devise a two-stage training pipeline to help the Spoken-LLM better learn the speaking styles. Based on extensive experiments, we show that Spoken-LLM outperforms text-only baselines and prior speech LLMs methods.  [^1] Demo of the StyleTalk dataset and output of Spoken-LLM are at https://sites.google.com/view/spoken-llm/home.
@@ 1 Introduction
Dataset S2S Expressive Purpose Diff styles&resp IEMOCAP recognition Switchboard recognition MUStARD recognition SEMAINE recognition MELD recognition MEISD recognition SCQA question answering NMSQA question answering $*{}^{*}$ question answering $*{}^{*}$ dialogue generation StyleTalk dialogue generation
Table 1: The list of spoken conversation datasets. “S2S" means speech-to-speech, and “Diff styles&resp" stands for the same sentence in different speaking styles and responses. In the “Purpose" column, “recognition" refers to recognizing the speaking style attributes in the speech, “question answering" means the task is formulated as the (question, answer) pair, and “dialogue generation" is the general chatbot agent to response any kinds of input. The datasets noted with $*{}^{*}$ are purely generated by LLM.
@@ 3.4 Spoken Dialogue Modeling
[…]
@@ 5.2 Warmup pre-training and data quality
5.3 Qualitative analysis
We show the qualitative example in Table 5 with different models’ outputs. This example shows that the Text-LLM (text-only) baseline predicts a more neutral sentiment response text, while the text-LLM (cascaded) and Spoken-LLM model generate text with a more aggressive and engaging tone, and the predicted response speaking styles are similar to the ground truth.
@@ 6 Related works
Modeling Speaking Style in Spoken Dialogue: Speaking style is important for speech understanding and response generation in spoken dialogue. The understanding of speaking styles in spoken dialogue is crucial for extracting style attributes such as emotion, sentiment, sarcasm, and more. Representative corpora for studying speaking styles in spoken conversations include IEMOCAP (Busso et al., 2008), SEMAINE (McKeown et al., 2010), MUStARD (Castro et al., 2019), Switchboard-sentiment (Chen et al., 2020), MELD (Poria et al., 2019), and MEISD (Firdaus et al., 2020). These datasets are primarily constructed based on label annotations from real speech conversations (e.g., TV series) or acted spoken conversations.
@@ 7 Conclusion
Direct speech-to-speech modeling: The Spoken-LLM generates predefined style attributes for inputting into the expressive TTS system. Future work on directly modeling responding speech has the potential to eliminate the need for explicit style labels.
@@
[…]
@@ Appendix I Dataset license
We plan to release the StyleTalk dataset under the MIT license.

### v2

@@ Abstract
In spoken dialogue, even if two current turns are the same sentence, their responses might still differ when they are spoken in different styles. The spoken styles, containing paralinguistic and prosodic information, mark the most significant difference between text and speech modality. When using text-only LLMs to model spoken dialogue, text-only LLMs cannot give different responses based on the speaking style of the current turn. In this paper, we focus on enabling LLMs to listen to the speaking styles and respond properly. Our goal is to teach the LLM that "even if the sentences are identical if they are spoken in different styles, their corresponding responses might be different". Since there is no suitable dataset for achieving this goal, we collect a speech-to-speech dataset, StyleTalk, with the following desired characteristics: when two current speeches have the same content but are spoken in different styles, their responses will be different. To teach LLMs to understand and respond properly to the speaking styles, we propose the Spoken-LLM framework that can model the linguistic content and the speaking styles. We train Spoken-LLM using the StyleTalk dataset and devise a two-stage training pipeline to help the Spoken-LLM better learn the speaking styles. Based on extensive experiments, we show that Spoken-LLM outperforms text-only baselines and prior speech LLMs methods.  [^1] Demo of the StyleTalk dataset and output of Spoken-LLM are at https://sites.google.com/view/spoken-llm/home. Code and dataset are available at https://github.com/DanielLin94144/StyleTalk.
@@ 1 Introduction
Dataset S2S Expressive Purpose Diff styles&resp IEMOCAP recognition Switchboard recognition MUStARD recognition SEMAINE recognition MELD recognition MEISD recognition MSP-improv recognition SCQA question answering NMSQA question answering OpenSAQA question answering E-chat200 dialogue generation StyleTalk dialogue generation
Table 1: The list of spoken conversation datasets. “S2S" means speech-to-speech, and “Diff styles&resp" stands for the same sentence in different speaking styles and responses. In the “Purpose" column, “recognition" refers to recognizing the speaking style attributes in the speech, “question answering" means the task is formulated as the (question, answer) pair, and “dialogue generation" is the general chatbot agent to response any kinds of input. The datasets noted with are purely generated by LLM.
@@ 3.4 Spoken Dialogue Modeling
Method Response text Response style BLEU ROUGEl METEOR BERTf1 F1emotion F1speed F1volume Text-LLM (text-only) 3.1 16.2 17.4 75.3 17.5 37.1 41.9 Text-LLM (cascaded) 3.2 17.3 19.1 76.0 37.5 52.9 65.6 Text-LLM (upper bound) 4.0 17.9 19.6 76.3 40.2 53.5 65.8 ParalinGPT-utt 3.1 16.8 18.5 75.9 32.3 51.9 64.8 ParalinGPT-chunk 3.1 16.5 18.2 75.8 34.0 54.8 65.8 Spoken-LLM-utt 2.8 16.6 20.2 75.8 47.4 61.5 56.5 Spoken-LLM-chunk 4.0 17.8 19.4 76.3 49.6 62.1 61.1
@@ 5.2 Warmup pre-training and data quality
5.3 Qualitative example
We show the qualitative example in Table 5 with different models’ outputs. This example shows that the Text-LLM (text-only) baseline predicts a more neutral sentiment response text, while the text-LLM (cascaded) and Spoken-LLM model generate text with a more aggressive and engaging tone, and the predicted response speaking styles are similar to the ground truth. More quantitative analyses are in Appendix D (style transition) and E (Diversity of current and responding style).
@@ 6 Related works
Speaking Style in Spoken Dialogue: Speaking style is important for speech understanding and response generation in spoken dialogue. The understanding of speaking styles in spoken dialogue is crucial for extracting style attributes such as emotion, sentiment, sarcasm, and more. Representative corpora for studying speaking styles in spoken conversations include IEMOCAP (Busso et al., 2008), SEMAINE (McKeown et al., 2010), MUStARD (Castro et al., 2019), Switchboard-sentiment (Chen et al., 2020), MELD (Poria et al., 2019), MEISD (Firdaus et al., 2020), and MSP-improv (Busso et al., 2016). These datasets are primarily constructed based on label annotations from real speech conversations (e.g., TV series) or acted spoken conversations.
@@ 7 Conclusion
Direct speech-to-speech modeling: The Spoken-LLM generates predefined style attributes to feed into the expressive TTS system. Future work on directly modeling responding speech has the potential to eliminate the need for explicit style labels.
@@
Acknowledgement
We want to express our gratitude to the reviewers for their detailed feedback and actionable suggestions, which helped us strengthen our paper. Guan-Ting Lin is supported by a joint NTU GICE and Nvidia Ph.D. scholarship program. Additionally, we thank the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.
@@ Acknowledgement
Method Response text Response style BLEU ROUGEl METEOR BERTf1 F1emotion F1speed F1volume Text-LLM (cascaded) 3.1 (-0.1) 16.9 (-0.4) 18.5 (-0.6) 75.9 (-0.1) 37.0 (-0.5) 52.5 (-0.4) 63.7 (-2.1) Spoken-LLM-chunk 3.3 (-0.7) 17.1 (-0.9) 19.0 (-0.4) 75.9 (-0.4) 47.5 (-2.1) 60.3 (-1.8) 57.4 (-3.7)
@@ Appendix I Dataset license
We released the StyleTalk dataset under the MIT license.